OpenStudy (anonymous):

Which is a factor of x^6 – 27y^3 ?

OpenStudy (anonymous):

Use the formula for a difference in cubes. \[a^3 - b^3 = (a - b)(a^2 + ab + b^2)\]

OpenStudy (anonymous):

okay .

OpenStudy (anonymous):

what does a b and c equal though ?

OpenStudy (anonymous):

There is no c? a = x^2 and b = 3y

OpenStudy (anonymous):

oh duh ! okay .

OpenStudy (anonymous):

Got it from here?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Alright. Tell me what you get.

OpenStudy (anonymous):

okay .